Find the slope of the line through each pair of points
( -1, 17), (-8, 16)
(-2, -20), (-18, -7)

Answer by stanbon(75887)

Find the slope of the line through each pair of points
----
slope = change in y/change in x
------
( -1, 17), (-8, 16)
slope = (16-17)/(-8--1) = -1/-7 = 1/7
-------------------------
(-2, -20), (-18, -7)
slope = (-7--20)/(-18--2) = 13/-16 = -13/16
-------------
Cheers,
Stan H.
------------------